Determine the pH at the equivalence (stoichiometric) point in the titration of 35 mL of 0.12 M H3BO3(aq) with 0.25 M NaOH(aq). The Ka of H3BO3 is 5.8 x 10-10. Ans: 11.07
Answer #1

at equivalence Miv, = point M₂ V₂ M₂=0.25 m V₂ = ? Mi = 0.12 M = 3 sme 0-12x35 z = 0.25 - 16.8 total volume = 16.8 +35 = 51.8